After robbing a bank in Dodge City, a robber gallops off at 12 mi/h. 10 minutes later, the marshall leaves to pursue the robber at 15 mi/h. How long (in hours) does it take the marshall to catch up to the robber?

Answer by checkley77(12844)

D=RT
D=12*T FOR THE BANK ROBBER.
D=15(T-1/6)
12T=15T-15/6
15T-12T=5/2
3T=2.5
T=2.5/3
T=.8333 HOURS TO CATCH THE ROBBER.
PROOF:
12*.8333=15(.8333-.1666)
10=15*.6667
10=10